# Empirical Networks are Sparse: Enhancing Multi-Edge Models with Zero-Inflation

### v1

@@ Abstract
Real-world networks are sparse. As we show in this article, even when a large number of interactions is observed most node pairs remain disconnected. We demonstrate that classical multi-edge network models, such as the $G(N,p)$, configuration models, and stochastic block models, fail to accurately capture this phenomenon. To mitigate this issue, zero-inflation must be integrated into these traditional models. Through zero-inflation, we incorporate a mechanism that accounts for the excess number of zeroes (disconnected pairs) observed in empirical data. By performing an analysis on all the datasets from the Sociopatterns repository, we illustrate how zero-inflated models more accurately reflect the sparsity and heavy-tailed edge count distributions observed in empirical data. Our findings underscore that failing to account for these ubiquitous properties in real-world networks inadvertently leads to biased models which do not accurately represent complex systems and their dynamics.
@@ 1 Introduction
Consider the interactions between students in a high-school. Students from different classes have limited opportunities to meet due to distinct schedules and social circles. When they don’t share common spaces or schedules, they don’t interact. However, when they do meet, the frequency of their interactions can vary independently of their initial chance of meeting. This dynamic, characteristic of many real-world networks, often results in a phenomenon known as “zero-inflation” in the distribution of multi-edges. The number of observed zeroes—i.e., the number of disconnected node pairs—exceeds what we would expect, given the large number of potential interactions, as illustrated in Fig. 1(a).
In other words, empirical networks are typically sparse [16, 12, 25]: only a limited number of node pairs have multiple interactions (multi-edges), while the majority remains disconnected. Traditional network models, such as the $G(N,p)$ [15, 20], configuration models [8, 18, 6], and stochastic block models [23, 33], have been instrumental in advancing our understanding of complex networks. However, these models fall short in representing the inherent sparsity observed in real multi-edge network data. They usually assume a proportional relationship between the growth of edges and the number of connected node pairs, which is not always accurate.
@@
Figure 1: Empirical multi-edge networks are sparse. Traditional multi-edge models like the G(N,p) struggle to reflect real-world data characteristics. (a) edge count distribution in Zachary’s Karate Club showing bimodality. The red solid line represents the G(N,p) prediction and the blue dashed line its zero-inflated counterpart. (b) In grey, the number of multi-edges per pair $m/\binom{N}{2}$ and the fraction of connected node pairs $M/\binom{N}{2}$ over time, averaged over all Sociopatterns datasets. While $m$ rises sharply, $M$ exhibits a steadier growth, indicating that most interactions occur among existing pairs. The dashed red line denotes the expected fraction of connected pairs according to the G(N,p) model defined by the corresponding $m$ value. Note that while the model quickly predicts a fully connected network, the empirical network remains sparse.
@@ 1.2 Modelling zero-inflation
In addition to zero-inflated models, hurdle models provide another approach to address count data with excess zeroes [17]. While zero-inflated models combine a binary process for zero counts with a Poisson process for positive counts, hurdle models treat zeros and positive counts as outcomes of two distinct processes. Specifically, a hurdle model first uses a binary process to determine whether an interaction occurs (i.e., whether the count is zero or positive), and then applies a truncated Poisson process for positive counts. This separation ensures that zeros are generated differently than positive counts, which can offer advantages in terms of model identifiability and interpretability [29]. However, this assumption implies that sparsity is only generated by the hurdle process, and not by low interaction rates. Such a strict assumption may not always hold true, especially in cases where modelling the interplay between low interaction rates and zero-inflation is critical [17].
@@
Throughout this article, we consistently employ a variation on the method of moments for parameter inference. This approach involves setting the first moments of the respective models to observed quantities—such as the number of edges, number of links, and degree sequences—from a given network. This consistency in methodology facilitates a coherent comparison across different models and allows for a clear understanding of their individual and comparative characteristics.
Often, though, we would like to obtain the maximum likelihood estimate (MLE) of the model parameters. Unfortunately, MLE presents significant challenges for zero-inflated models, a complexity underscored by the existing literature [34, 2, 9]. The inherent difficulty arises particularly in the context of network models, given the increased dimensionality. The primary reason for this complexity is that a maximum likelihood estimate of the mixture probability in Eq. 4 involves a logarithm of sums, a form that does not lend itself to simplification or easy manipulation with standard tools. Consequently, optimising the log-likelihood function usually requires a numerical approach. In these cases, the estimates obtained by the methods-of-moments serve as initial values for the numerical MLE optimisation.
@@ 2.1 Zero-Inflated $G(N,p)$ Model (zi-$G(N,p)$)
To estimate the parameters $p$ and $q$ from observed data, we use the method of moments, matching these expected values to the observed counts of interactions and links, $\widehat{m}$ and $\widehat{M}$, respectively:
@@
$\widehat{p}=\frac{\widehat{m}}{N^{2}q}\,,\qquad\widehat{q}=\frac{\widehat{m}\widehat{M}}{N^{2}(\widehat{m}+\widehat{M}\mathcal{W}\left[-\frac{\widehat{m}e^{-\frac{\widehat{m}}{\widehat{M}}}}{\widehat{M}}\right])}\,,$ (10)
@@ 2.2 Zero-Inflated Stochastic Block Model (zi-SBM)
Building on the $G(N,p)$ model, the Stochastic Block Model (SBM) introduces a richer representation of network structures by distinguishing distinct blocks in the network, each characterised by a unique edge probability [22, 1]. This model captures community structures within networks. Nodes are assigned to one of $B$ different groups, and the probability of interactions between nodes $i$ and $j$ in groups $b_{i}$ and $b_{j}$, respectively, is determined by a block-specific parameter $\lambda_{b_{i}b_{j}}$.
@@
$\Pr(\mathcal{G}|\boldsymbol{\lambda},\boldsymbol{q})=\prod_{ij}\left((1-q_{b_{i}b_{j}})\delta_{0}(A_{ij})+q_{b_{i}b_{j}}\frac{\lambda_{b_{i}b_{j}}^{A_{ij}}}{A_{ij}!}\exp(-\lambda_{b_{i}b_{j}})\right),$ (11)
@@
$\mathbb{E}(M|\boldsymbol{\lambda},\boldsymbol{q})$ $=\sum_{bd}N_{b}N_{d}q_{bd}(1-e^{-\lambda_{bd}}),$ (13)
@@
A closed-form solution for $\widehat{q}_{bd}$ is not readily available. Nonetheless, the values of $\widehat{q}_{bd}$ can be determined by numerically solving the set of $B^{2}$ independent equations given in Eq. 14.
@@ 2.3 Zero-Inflated Configuration Model (zi-CLCM)
While both the $G(N,p)$ and SBM models offer valuable insights, they fall short in encoding node heterogeneities. Configuration models fill this gap by introducing a parameterisation that accounts for degree heterogeneities [18]. In the framework of Poisson models, the Chung-Lu configuration model (CLCM) achieves this by expressing the general parameters $\lambda_{ij}$ as $\theta^{\text{out}}_{i}\theta^{\text{in}}_{j}$, a product node-parameters [8, 31]. For undirected networks, we have $\boldsymbol{\theta^{\text{out}}}=\boldsymbol{\theta^{\text{in}}}=\boldsymbol{\theta}$.
@@
$\Pr(\mathcal{G}|\boldsymbol{\theta},q)=\prod_{ij}\left((1-q)\,\delta_{0}(A_{ij})+q\,\frac{(\theta^{\text{out}}_{i}\theta^{\text{in}}_{j})^{A_{ij}}}{A_{ij}!}\exp(-\theta^{\text{out}}_{i}\theta^{\text{in}}_{j})\right).$ (16)
It is important to note that the model parameters appear always in pairs. This means that they are defined modulo constants. This characteristic requires fixing a constraint that ensures a unique solution to the inference problem. A common approach is to fix the expected number of interactions in the network:
$\mathbb{E}(m|\boldsymbol{\theta},q)=q\,\sum_{ij}\theta^{\text{out}}_{i}\theta^{\text{in}}_{j}\,.$ (17)
To infer the parameters of the network model $q$, $\boldsymbol{\theta^{\text{out}}}$, and $\boldsymbol{\theta^{\text{in}}}$, we set four different moments to observed values:
$\mathbb{E}(m|\boldsymbol{\theta},q):=\widehat{m},\quad\mathbb{E}(M|\boldsymbol{\theta},q):=\widehat{M},\quad\forall i:\;\mathbb{E}(k^{\text{out}}_{i}|\boldsymbol{\theta},q):=\widehat{k^{\text{out}}_{i}},\quad\forall i:\;\mathbb{E}(k^{\text{in}}_{i}|\boldsymbol{\theta},q):=\widehat{k^{\text{in}}_{i}},$ (18)
where $\widehat{m}$ denotes the number of interactions in an observed network, $\widehat{M}$ represents its number of links, and $\widehat{k^{\text{out}}_{i}}$ and $\widehat{k^{\text{in}}_{i}}$ the observed out- and in-degrees of $i$.
The expected degree sequences and the number of links from the model are given by:
$\mathbb{E}(k^{\text{out}}_{i}|\boldsymbol{\theta},q)=q\,\theta^{\text{out}}_{i}\sum_{j}\theta^{\text{in}}_{j},\qquad\mathbb{E}(k^{\text{in}}_{i}|\boldsymbol{\theta},q)=q\,\theta^{\text{in}}_{i}\sum_{j}\theta^{\text{out}}_{j},$ (19)
$\mathbb{E}(M|\boldsymbol{\theta},q)=q\,N^{2}-q\,\sum_{ij}e^{-\theta^{\text{out}}_{i}\theta^{\text{in}}_{j}}\,.$ (20)
Fixing the expected number of interactions to $\widehat{m}$, we obtain a constraint for the L1-norm of the parameter vectors $\boldsymbol{\theta^{\text{out}}}$ and $\boldsymbol{\theta^{\text{in}}}$:
$\sum_{i}\theta^{\star}_{i}=\sqrt{\frac{\widehat{m}}{q}}.$ (21)
This constraint ties the expected number of interactions in the network to the node-specific parameters, allowing their inference.
From Eq. 21 and Eq. 19, we derive expressions for $\widehat{\theta^{\star}_{i}}$ as functions of $q$ and the observed degrees $\widehat{k^{\text{out}}_{i}}$ and $\widehat{k^{\text{in}}_{i}}$:
@@
Lastly, substituting Eq. 22 into Eq. 20 leads to an explicit equation for $\widehat{q}$:
$\widehat{q}\,\left(N^{2}-\sum_{ij}e^{-\frac{\widehat{k^{\text{out}}_{i}}\widehat{k^{\text{in}}_{j}}}{\widehat{m}\widehat{q}}}\right)=\widehat{M}.$ (23)
Solving numerically for $\widehat{q}$ allows the estimation of all model parameters.
@@ 2.4 Zero-Inflated Degree-Corrected Stochastic Block Model (zi-DCSBM)
We finally discuss the Zero-Inflated Degree Corrected Stochastic Block Model (zi-DCSBM). As for the standard degree-corrected stochastic block model [23], it incorporates both node heterogeneity and block structure into a single comprehensive model, combining the features of the zi-CLCM and the zi-SBM. The introduction of block-specific zero-inflation parameters $q_{bd}$, constituting the vector $\boldsymbol{q}$, allows the model to account for varying levels of zero-inflation across different blocks. By parametrising Eq. 2 with $\lambda_{ij}=\theta^{\text{out}}_{i}\theta^{\text{in}}_{j}\lambda_{b_{i}b_{j}}$, we derive the following probability distribution:
@@
$\Pr(\mathcal{G}|\boldsymbol{\theta},\boldsymbol{\lambda},\boldsymbol{q})=\prod_{ij}\left((1-q_{b_{i}b_{j}})\,\delta_{0}(A_{ij})+q_{b_{i}b_{j}}\,\frac{(\theta^{\text{out}}_{i}\theta^{\text{in}}_{j}\lambda_{b_{i}b_{j}})^{A_{ij}}}{A_{ij}!}\exp(-\theta^{\text{out}}_{i}\theta^{\text{in}}_{j}\lambda_{b_{i}b_{j}})\right).$ (24)
In order to infer the model parameters, we employ the method of moments, setting the first moments of the model to their observed counterparts in a given network. The expected values of the number of interactions, node degrees, and links for each block are defined as follows:
$\mathbb{E}(m_{bd}|\boldsymbol{\theta},\boldsymbol{\lambda},\boldsymbol{q})=q_{bd}\lambda_{bd}\sum_{i\in b}\theta^{\text{out}}_{i}\sum_{j\in d}\theta^{\text{in}}_{j},$ (25)
$\mathbb{E}(k^{\text{out}}_{i}|\boldsymbol{\theta},\boldsymbol{q})=\theta^{\text{out}}_{i}\sum_{d}q_{b_{i}d}\lambda_{b_{i}d}\sum_{j\in d}\theta^{\text{in}}_{j},\qquad\mathbb{E}(k^{\text{in}}_{i}|\boldsymbol{\theta},\boldsymbol{q})=\theta^{\text{in}}_{i}\sum_{d}q_{b_{i}d}\lambda_{b_{i}d}\sum_{j\in d}\theta^{\text{out}}_{j},$ (26)
$\mathbb{E}(M_{bd}|\boldsymbol{\theta},\boldsymbol{\lambda},\boldsymbol{q})=q_{bd}\left(N_{b}N_{d}-\sum_{i\in b,j\in d}e^{-\lambda_{bd}\theta^{\text{out}}_{i}\theta^{\text{in}}_{j}}\right),$ (27)
where $N_{b}$ denotes the number of nodes in group $b$.
Just like in the CLCM and zi-CLCM, the parameters $\boldsymbol{\theta^{\text{out}}}$ and $\boldsymbol{\theta^{\text{in}}}$ in the DCSBM and zi-DCSBM are defined modulo constants, necessitating the establishment of a constraint to ensure a unique solution to the inference problem. A common approach is to constrain the L1-norm of the node-specific parameters, i.e., $\sum_{i}\theta^{\star}_{i}\delta_{b}(b_{i})=1$. By incorporating this constraint into Eqs. 25 and 26, we obtain:
$\mathbb{E}(m_{bd}|\boldsymbol{\theta},\boldsymbol{\lambda},\boldsymbol{q})=q_{bd}\lambda_{bd},\quad\mathbb{E}(k^{\star}_{i}|\boldsymbol{\theta},\boldsymbol{q})=\theta^{\star}_{i}\sum_{d}q_{b_{i}d}\lambda_{b_{i}d}.$ (28)
@@
Solving the set of equations defined by Eqs. 25 and 26 for $\boldsymbol{\lambda}$, $\boldsymbol{\theta^{\text{out}}}$, and $\boldsymbol{\theta^{\text{in}}}$ yields expressions for their estimators that depend on the mixture parameters $\boldsymbol{q}$ and observed quantities:
$\widehat{\lambda}_{bd}=\frac{\widehat{m}_{bd}}{q_{bd}},\quad\theta^{\star}_{i}=\frac{\widehat{k^{\star}_{i}}}{\widehat{\kappa}^{\star}_{b_{i}}},$ (29)
where $\kappa^{\text{out}}_{b_{i}}=\sum_{d}m_{b_{i}d}$ and $\kappa^{\text{in}}_{b_{i}}=\sum_{d}m_{db_{i}}$ denote the out- and in-degree of block $b_{i}$, respectively.
Finally, substituting these expressions into Eq. 27 provides a set of equations for each $q_{bd}$ and the number of links in each block $\widehat{M}_{bd}$:
$q_{bd}\,\left(N_{b}N_{d}-\sum_{i\in b,j\in d}e^{-\frac{\widehat{k^{\text{out}}_{i}}\widehat{k^{\text{in}}_{j}}}{\widehat{\kappa}^{\text{out}}_{b_{i}}\widehat{\kappa}^{\text{in}}_{b_{j}}}\frac{\widehat{m}_{b_{i}b_{j}}}{q_{bd}}}\right)=\widehat{M}_{bd}.$ (30)
Numerically solving this set of $B^{2}$ independent equations yields estimates for all model parameters.
@@
$\Pr(\mathcal{G}|\boldsymbol{\theta},\boldsymbol{\lambda},\boldsymbol{q})=\prod_{ij}\begin{cases}(1-q_{b_{i}b_{j}})\quad\text{if }A_{ij}=0\,,\\ q_{b_{i}b_{j}}\,\frac{(\theta^{\text{out}}_{i}\theta^{\text{in}}_{j}\lambda_{b_{i}b_{j}})^{A_{ij}}\exp(-\theta^{\text{out}}_{i}\theta^{\text{in}}_{j}\lambda_{b_{i}b_{j}})}{A_{ij}!\left(1-\exp(-\theta^{\text{out}}_{i}\theta^{\text{in}}_{j}\lambda_{b_{i}b_{j}})\right)}\quad\text{if }A_{ij}>0\,,\end{cases}$ (31)

### v2

@@ Abstract
Real-world networks are sparse. As we show in this article, even when a large number of interactions is observed, most node pairs remain disconnected. We demonstrate that classical multi-edge network models, such as the $G(N,p)$, configuration models, and stochastic block models, fail to accurately capture this phenomenon. To mitigate this issue, zero-inflation must be integrated into these traditional models. Through zero-inflation, we incorporate a mechanism that accounts for the excess number of zeroes (disconnected pairs) observed in empirical data. By performing an analysis on all the datasets from the Sociopatterns repository, we illustrate how zero-inflated models more accurately reflect the sparsity and heavy-tailed edge count distributions observed in empirical data. Our findings underscore that failing to account for these ubiquitous properties in real-world networks inadvertently leads to biased models that do not accurately represent complex systems and their dynamics.
@@ 1 Introduction
Consider the interactions between students in a high school. Interactions can be represented as edges in a network where students are the nodes. Pairs of students who do not interact appear as disconnected, while repeated interactions between the same pair of students create multi-edges (also known as parallel edges), where the multiplicity of each edge represents the interaction count. Students from different classes have limited opportunities to meet due to distinct schedules and social circles. When they don’t share common spaces or schedules, they don’t interact. However, when they do meet, the frequency of their interactions can vary independently of their initial chance of meeting.
This dynamic, common in real-world networks, often leads to “zero-inflation” in the distribution of interaction counts: the number of disconnected pairs exceeds what we would expect given the large number of potential interactions. This results in an inflated number of zeroes in the distribution of interaction counts, as illustrated in Fig. 1 (A).
In other words, empirical networks are typically sparse [18, 13, 28]: only a limited number of node pairs share multiple interactions, while the majority remains disconnected. Traditional network models, such as the $G(N,p)$ [17, 22], configuration models [9, 20, 7], and stochastic block models [26, 36], have been instrumental in advancing our understanding of complex networks. However, these models fall short in representing the inherent sparsity observed in real multi-edge network data. They usually assume a proportional relationship between the growth of edges and the number of connected node pairs, which is not always accurate.
@@
Figure 1: Empirical multi-edge networks are sparse. Traditional multi-edge models like the G(N,p) struggle to reflect real-world data characteristics. (A) Edge count distribution in Zachary’s Karate Club showing bimodality. The red solid line represents the G(N,p) prediction, and the blue dashed line its zero-inflated counterpart. (B) Over time, edges accumulate between the same pairs of nodes in real-world networks. In grey, the interquartile range of the number of multi-edges per pair $\rho=m/\binom{N}{2}$ and the fraction of connected node pairs $d=M/\binom{N}{2}$, over all Sociopatterns datasets. The black lines denote the median values, while the dashed red line represents the expected fraction of connected pairs according to the G(N,p) model with the corresponding $m$ value. Note that while the model quickly predicts a fully connected network, the empirical network remains sparse, indicating that most interactions occur among existing pairs rather than forming new connections.
@@ 1.2 Modelling zero-inflation
In addition to zero-inflated models, hurdle models provide another approach to address count data with excess zeroes [19]. While zero-inflated models combine a binary process for zero counts with a Poisson process for positive counts, hurdle models treat zeros and positive counts as outcomes of two distinct processes. Specifically, a hurdle model first uses a binary process to determine whether an interaction occurs (i.e., whether the count is zero or positive), and then applies a truncated Poisson process for positive counts. This separation ensures that zeros are generated differently than positive counts, which can offer advantages in terms of model identifiability and interpretability [32]. However, this assumption implies that sparsity is only generated by the hurdle process, and not by low interaction rates. Such a strict assumption may not always hold true, especially in cases where modelling the interplay between low interaction rates and zero-inflation is critical [19]. Finally, if the hurdle is set at zero, as most commonly done, it can be shown that the resulting hurdle model is just a re-parametrized zero-inflated model [25].
@@
Throughout this article, we employ a variation of the Expectation-Maximization (EM) algorithm for parameter inference, specifically tailored to leverage key properties of zero-inflated Poisson (ZIP) distributions. This approach ensures methodological consistency across different models, allowing for coherent comparisons and a clearer understanding of their individual characteristics.
Direct Maximum Likelihood Estimation (MLE) for zero-inflated models presents significant challenges, as noted in prior work [37, 2, 10]. The difficulty stems from the complexity of network models in high-dimensional spaces, where the MLE of the mixture probability in Eq. 4 involves a logarithm of sums. This form complicates simplification and resists manipulation with standard tools, making numerical optimization of the log-likelihood function necessary.
Rather than directly estimating the parameters $q$ and $\lambda$ in Eq. 4, we leverage the fact that the ZIP distribution is a power series distribution (PSD). As shown in [25], PSDs allow us to use the first moment equation as a maximum-likelihood equation. Solving
$\bar{x}=\mathbb{E}[n]=\widehat{\lambda}\widehat{q}$ (5)
yields one of the maximum-likelihood equations for parameter estimation.
In the univariate case, the second MLE equation is defined by matching the observed proportion of zeroes $f_{0}$ to its expected value [25]:
$(1-\widehat{q})+\widehat{q}e^{-\widehat{\lambda}}=f_{0}.$ (6)
In the univariate case of Eq. 4, fixing the first two moments–one for the count process and one for the zero process–effectively achieves the MLE of the parameters. For the more complex models discussed below, however, the second equation may not always be applicable due to dependencies between parameters. In network models with structured dependencies (e.g., degree-corrected models), the mixture parameter $q$ may vary by node pair or community, introducing correlations that complicate isolating the zero-inflation component.
To address this, we rely on the first equation, which remains valid as an MLE condition, to link the Poisson and mixture parameters. We then treat the Poisson parameters as latent and optimize over the mixture ones, effectively defining a one-step EM algorithm.
@@ 2.1 Zero-Inflated $G(N,p)$ Model (zi-$G(N,p)$)
As the zi-$G(N,p)$ model corresponds to the univariate ZIP distribution discussed above, the maximum likelihood estimates of the parameters $p$ and $q$ can be obtained by matching these expected values to the observed counts of interactions and links, $\widehat{m}$ and $\widehat{M}$, respectively:
@@
$\widehat{p}=\frac{\widehat{m}}{N^{2}\widehat{q}}\,,\qquad\widehat{q}=\frac{\widehat{m}\widehat{M}}{N^{2}(\widehat{m}+\widehat{M}\mathcal{W}\left[-\frac{\widehat{m}e^{-\frac{\widehat{m}}{\widehat{M}}}}{\widehat{M}}\right])}\,,$ (12)
@@ 2.2 Zero-Inflated Stochastic Block Model (zi-SBM)
Building on the $G(N,p)$ model, the Stochastic Block Model (SBM) introduces a richer representation of network structures by distinguishing blocks in the network, each characterised by a unique edge probability [24, 1]. By doing so, the model captures community structures. Nodes are assigned to one of $B$ different groups, and the probability of interactions between nodes $i$ and $j$ in groups $b_{i}$ and $b_{j}$, respectively, is determined by a block-specific parameter $\lambda_{b_{i}b_{j}}$. Hereafter, we assume block assignments to be known a priori. We discuss the implications for community detection of zero-inflation later in the manuscript.
@@
$\begin{split}\Pr&(\mathcal{G}|\boldsymbol{\lambda},\boldsymbol{q})=\\ &\prod_{ij}\left((1-q_{b_{i}b_{j}})\delta_{0}(A_{ij})+q_{b_{i}b_{j}}\frac{\lambda_{b_{i}b_{j}}^{A_{ij}}}{A_{ij}!}\exp(-\lambda_{b_{i}b_{j}})\right),\end{split}$ (13)
@@
$\mathbb{E}(M_{bd}|\boldsymbol{\lambda},\boldsymbol{q})=\sum_{bd}N_{b}N_{d}q_{bd}(1-e^{-\lambda_{bd}}),$ (15)
@@
A closed-form solution for $\widehat{q}_{bd}$ is not readily available. Nonetheless, the values of $\widehat{q}_{bd}$ can be determined by numerically solving the set of $B^{2}$ independent equations given in Eq. 16. Because the zi-SBM effectively defines B independent zi-$G(N,p)$ models, i.e., one per block, $\widehat{\lambda}_{bd}$ and $\widehat{q}_{bd}$ are the maximum likelihood estimates for the model parameters.
@@ 2.3 Zero-Inflated Configuration Model (zi-CLCM)
While both the $G(N,p)$ and SBM models offer valuable insights, they fall short in encoding node heterogeneities. Configuration models fill this gap by introducing a parameterisation that accounts for degree heterogeneities [20]. In the framework of Poisson models, the Chung-Lu configuration model (CLCM) achieves this by expressing the general parameters $\lambda_{ij}$ as $\theta^{\text{out}}_{i}\theta^{\text{in}}_{j}$, a product of node-parameters [9, 34]. For undirected networks, we have $\boldsymbol{\theta^{\text{out}}}=\boldsymbol{\theta^{\text{in}}}=\boldsymbol{\theta}$.
@@
$\begin{split}\Pr&(\mathcal{G}|\boldsymbol{\theta},q)=\\ &\prod_{ij}\left((1-q)\,\delta_{0}(A_{ij})+q\,\frac{(\theta^{\text{out}}_{i}\theta^{\text{in}}_{j})^{A_{ij}}}{A_{ij}!}\exp(-\theta^{\text{out}}_{i}\theta^{\text{in}}_{j})\right).\end{split}$ (18)
Parameters Identifiability
It is important to note that the model’s Poisson parameters $\boldsymbol{\theta^{\text{out}}}$ and $\boldsymbol{\theta^{\text{in}}}$ appear always in pairs. This means that they are defined modulo constants. This characteristic requires fixing a constraint to enable their estimation. A common approach is to re-parametrize the model by introducing a new parameter $C$ that constrains the L1-norm of $\boldsymbol{\theta^{\text{out}}}$ and $\boldsymbol{\theta^{\text{in}}}$:
$\sum_{i}\theta^{\star}_{i}=C,$ (19)
where the superscript stands for in/out.
Parameters Estimation
To define the first set of maximum likelihood equations, corresponding to Eq. 5 for the univariate case, we set three different first-moment equations to their observed values:
$\mathbb{E}(m|\boldsymbol{\theta},q):=\widehat{m},$ (20)
$\forall i:\;\mathbb{E}(k^{\text{out}}_{i}|\boldsymbol{\theta},q):=\widehat{k^{\text{out}}_{i}},$ $\quad\forall i:\;\mathbb{E}(k^{\text{in}}_{i}|\boldsymbol{\theta},q):=\widehat{k^{\text{in}}_{i}},$ (21)
where $\widehat{m}$ denotes the number of interactions in an observed network and $\widehat{k^{\text{out}}_{i}}$ and $\widehat{k^{\text{in}}_{i}}$ the observed out- and in-degrees of $i$.
Replacing Eq. 19 in Eq. 20 yields
$\mathbb{E}(m|\boldsymbol{\theta},q)=q\sum_{i,j}\theta^{\text{out}}_{i}\theta^{\text{in}}_{j}=qC^{2},$ (22)
which allows to tie the maximum-likelihood estimator $\widehat{C}$ to $q$:
$\widehat{C}=\sqrt{\frac{\widehat{m}}{q}}\,.$ (23)
The expected degree sequences from the model are given by:
$\begin{split}\mathbb{E}(k^{\text{out}}_{i}|\boldsymbol{\theta},q)=q\,\theta^{\text{out}}_{i}\sum_{j}\theta^{\text{in}}_{j},\\ \mathbb{E}(k^{\text{in}}_{i}|\boldsymbol{\theta},q)=q\,\theta^{\text{in}}_{i}\sum_{j}\theta^{\text{out}}_{j},\end{split}$ (24)
From Eqs. 19, 24 and 23, we derive expressions for $\widehat{\theta^{\star}_{i}}$ as functions of $q$ and the observed degrees $\widehat{k^{\text{out}}_{i}}$ and $\widehat{k^{\text{in}}_{i}}$:
@@
Lastly, we need to estimate the remaining parameter $q$. In this case, however, we need to resort to the explicit maximization of the likelihood function. Substituting Eq. 25 into Eq. 18 leads to an equation for $\widehat{q}$ that can be optimized numerically.
@@ 2.4 Zero-Inflated Degree-Corrected Stochastic Block Model (zi-DCSBM)
Finally, we discuss the Zero-Inflated Degree Corrected Stochastic Block Model (zi-DCSBM). As for the standard degree-corrected stochastic block model [26], it incorporates both node heterogeneity and block structure into a single comprehensive model, combining the features of the zi-CLCM and the zi-SBM. The introduction of block-specific mixture parameters $q_{bd}$, constituting the vector $\boldsymbol{q}$, allows the model to account for varying levels of zero-inflation across different blocks. By parametrising Eq. 2 with $\lambda_{ij}=\theta^{\text{out}}_{i}\theta^{\text{in}}_{j}\lambda_{b_{i}b_{j}}$, we derive the following probability distribution:
@@
$\begin{split}\Pr(\mathcal{G}|\boldsymbol{\theta},\boldsymbol{\lambda},\boldsymbol{q})&=\prod_{ij}\biggl{(}(1-q_{b_{i}b_{j}})\,\delta_{0}(A_{ij})+\\ &q_{b_{i}b_{j}}\,\frac{(\theta^{\text{out}}_{i}\theta^{\text{in}}_{j}\lambda_{b_{i}b_{j}})^{A_{ij}}}{A_{ij}!}\exp(-\theta^{\text{out}}_{i}\theta^{\text{in}}_{j}\lambda_{b_{i}b_{j}})\biggr{)}.\end{split}$ (26)
Parameters Identifiability
Just like in the CLCM and zi-CLCM, the parameters $\boldsymbol{\theta^{\text{out}}}$ and $\boldsymbol{\theta^{\text{in}}}$ in the DCSBM and zi-DCSBM are defined modulo constants. We re-parametrize the model by deriving $2B$ constraints for the L1-norm of the node-specific Poisson parameters:
$\sum_{i}\theta^{\star}_{i}\delta_{b}(b_{i})=C^{\star}_{b},$ (27)
for each block $b$. In contrast to the CLCM and zi-CLCM, here, the zi-DCSBM does not depend on the choice of parameters $C^{\star}_{b}$, as shown in Supplementary Information E. Therefore, we are free to set $C^{\star}_{b}:=C=1$ $\forall b$ to simplify the parameter estimation.
Parameters Estimation
Manipulating the first moment equation, we get the expected values of the number of interactions per block:
$\widehat{m}_{bd}=\mathbb{E}(m_{bd}|\boldsymbol{\theta},\boldsymbol{\lambda},\boldsymbol{q})=q_{bd}\lambda_{bd}\sum_{i\in b}\theta^{\text{out}}_{i}\sum_{j\in d}\theta^{\text{in}}_{j}.$ (28)
The constraints in Eq. 27 allow us to tie the block Poisson parameters $\boldsymbol{\lambda}$ to the mixture parameters $\boldsymbol{q}$:
$\widehat{\lambda}_{bd}=\frac{\widehat{m}_{bd}}{q_{bd}}.$ (29)
From the first moment equation, we derive the expected degree sequences
$\begin{split}\mathbb{E}(k^{\text{out}}_{i}|\boldsymbol{\theta},\boldsymbol{q})=\theta^{\text{out}}_{i}\sum_{d}q_{b_{i}d}\lambda_{b_{i}d}\sum_{j\in d}\theta^{\text{in}}_{j},\\ \mathbb{E}(k^{\text{in}}_{i}|\boldsymbol{\theta},\boldsymbol{q})=\theta^{\text{in}}_{i}\sum_{d}q_{b_{i}d}\lambda_{b_{i}d}\sum_{j\in d}\theta^{\text{out}}_{j}.\end{split}$ (30)
Leveraging the constraints given by Eq. 27, from Eq. 30, we obtain:
$\mathbb{E}(k^{\star}_{i}|\boldsymbol{\theta},\boldsymbol{q})=\theta^{\star}_{i}\sum_{d}q_{b_{i}d}\lambda_{b_{i}d}.$ (31)
@@
Solving the set of equations defined by Eqs. 29 and 31 for $\boldsymbol{\theta^{\text{out}}}$ and $\boldsymbol{\theta^{\text{in}}}$ yields their MLE:
$\widehat{\theta^{\star}_{i}}=\frac{\widehat{k^{\star}_{i}}}{\widehat{\kappa}^{\star}_{b_{i}}},$ (32)
where $\widehat{\kappa}^{\text{out}}_{b_{i}}=\sum_{d}\widehat{m}_{b_{i}d}$ and $\widehat{\kappa}^{\text{in}}_{b_{i}}=\sum_{d}\widehat{m}_{db_{i}}$ denote the out- and in-degree of block $b_{i}$, respectively. Finally, substituting these expressions into Eq. 26 provides a set of $B^{2}$ independent equations for each $q_{bd}$ that need to be optimized numerically to find the MLEs of all remaining parameters.
@@
$\begin{split}&\Pr(\mathcal{G}|\boldsymbol{\theta},\boldsymbol{\lambda},\boldsymbol{q})=\\ &\prod_{ij}\begin{cases}(1-q_{b_{i}b_{j}})\quad\text{if }A_{ij}=0\,,\\ q_{b_{i}b_{j}}\,\frac{(\theta^{\text{out}}_{i}\theta^{\text{in}}_{j}\lambda_{b_{i}b_{j}})^{A_{ij}}\exp(-\theta^{\text{out}}_{i}\theta^{\text{in}}_{j}\lambda_{b_{i}b_{j}})}{A_{ij}!\left(1-\exp(-\theta^{\text{out}}_{i}\theta^{\text{in}}_{j}\lambda_{b_{i}b_{j}})\right)}\quad\text{if }A_{ij}>0\,,\end{cases}\end{split}$ (33)